User Tools

Site Tools


lbaops:telecon2009jun11

VLBI teleconference, 11 june 2009

12.00 Perth, 14.00 Eastern, 16.00 New Zealand

Phone:

Vic:03 9545 8222
NSW:02 9490 8022
ACT:02 6276 6622
Qld:07 3214 2311
WA:08 6436 8844
SA:08 8273 8122
NT:08 8944 8488
Tas:03 6232 5188

Meeting ID: 2815

Chair: Tasso Tzioumis

Attendance:Oleg Titov (GA, Canberra); Shinji Horiuchi (SDCCC, Canberra);Claire Hotan, Aidan Hotan, Aquib Moin, Cormac Reynolds, Hayley Bignall (Curtin); Simon Ellingsen (UTas); Tzioumis, Phillips (ATNF Marsfield); Richard Dodson, Maria Rioja (UWA)

Apologies: Phil Edwards, Steven Tingay

Agenda

  • Observations in July 2009
    • General issues (Tzioumis/Phillips/Edwards)
    • Schedules (Edwards/C.Reynolds)
    • Support (Tzioumis)
    • Disks (Claire)
    • Mk5 disks (Chris/Tasso)
    • Transfers (Claire/Hayley/Chris)
  • REPORTS/UPDATES
  • Curtin cluster at Narrabri (Phillips/Amy)
  • Update on ASKAP VLBI EoI (Tingay)
  • Feb 2010 VLBI meetings
    • IVS meeting in Hobart (Lovell)
    • Astro meetings in NZ (Tingay)
    • LBA ops/eVLBI mtg?
  • Update on New Zealand antenna (Gulyaev)
  • AuScope Update (Lovell)
  • AOB

Actions arising

(new items in blue)

GENERAL

  • Swinburne archive of correlated VLBI data to be moved to ATNF & Curtin at end of all Swinburne correlation (Steven/Claire) - update 28/01/09 - Swinburne correlation complete. Claire transferring data to Curtin. Will tidy up before sending on to ATNF. Claire to advise ATNF on archive size. 11/2/09 - in progress. 21/5/09: All data in archive. Will check and make available to ATNF.Continuing (11/6/09)
    • Determine what meta-data should be archived (ATNF/Curtin)
    • Determine what muct be included in the user-accessible archive vs. archive at correlator (atnf/Curtin)
  • Provide additional advice to PIs on scheduling the ATCA to keep it phased up at 8 GHz and above (Tasso) - update 28/01/09 - Richard to have input regards polarisation properties. 11/2/09 - to be tested with observations in next session. Richard provided report. To be disseminated to community. (21/5/09)
    • New CABB systems may need experimentation to check best strategy and implementation for phasing the array. (Phil/Chris/Tasso)
  • List on the Wiki available bands/polarisations at each LBA antenna. Also update “standard bands” (from S2) and define “standard” bands for wide-band recording. (Cormac/Tasso). 11/2/09 - initial progress within SCHED. Will be further followed up.
    • Review what information is already available in freq.dat in SCHED (Chris/Cormac)
  • Curtin Cluster at Narrabri: Networking operational and configuration to be done on 12/06/09. To be brought to operation (Chris/Shaun/Jamie)
    • 36 TB store on cluster may need configuring. Steven initiated contact with IVEC expert who is working with Chris.
  • ARCS transfers.
    • Prepare checklist for procedures for next session. (Claire)
    • Get observers to update disk usage on wiki (all)
    • Prepare for 2nd phase of implementation of transfers, to include Hobart, Tid (via Canberra ARCS node) and Ceduna (via Adelaide ARCS node). May need XRAIDs at ARCS nodes. To be explored with ARCS contacts (Steven).
      • Hobart transfers via ARCS to be tried in July.
      • Explore doing Tid transfers vis USB/eSATA disk copy to ARCS in Canberra (Chris/Claire)
    • Communication re networks (Claire)
      • Transfer times - send explicit list of required downtimes to Graham
      • Networking requirements - nominate someone from each institution to liaise with Graham re: port requirements etc. and be available to make changes (If you'll be travelling, nominate a deputy)
  • Checklist for all LBA observations. Initial copy from Hobart (Brett). To be developed during next session (Tasso/Claire/Hayley)
    • Disk wiki to also show number of files at end of expt? (Claire)
  • Mk5 Disks. New units to be ordered for correlation in Curtin and/or Bonn. Steven and Tasso to coordinate.
  • Proposal to always correlate LBA data with all 4 Stokes, unless PI objects. Richard to write short report and justification, after checking with ATCA data.
  • Unified LBA monitoring and control, to include Ho & Cd. Initial tests done in May. Will be tried more fully in July. (Simon/Jamie/Tasso)

NEXT OBSERVING SESSION

  • Observing support for next session DONE
    • Observers organised for ATNF antennas. Tasso to coordinate.
    • Disks to observatories. Plenty of disks at Curtin. Claire/Tasso to coordinate
      • Disk list from Parkes to Claire (Ettore/Tasso)
      • Check disk inventory and ship. May need to ship only to Narrabri. (Claire)

OTHER

  • Collaborations with South America may be possible, to compensate for loss of Hart. TIGO (Chile), O'Higgins (Antarctica), Fortaleza (Brazil) & antenna in Peru. Tasso to investigate.
    • Small antennas to South America. Will continue investigating.
  • Next LBA ops/e-VLBI meeting next February, during IVS week in Hobart (8-12th). May be able to have LBA meeting Thu-Fri (11-12). Tasso to explore and coordinate.
  • Possible APAN e-VLBI demo in February, same period as IVS meeting. To explore further and minimise conflict. (Chris/Tasso)
  • Next telecon: Thu, 29 July 2009, 3pm Sydney time (Tasso/Chris/Cormac)
lbaops/telecon2009jun11.txt · Last modified: 2015/12/18 16:39 (external edit)